OpenStudy (caboose2):

simplify and write the answer in standard form (-5mn^3)(4m^2n^2)

OpenStudy (campbell_st):

the law for multiplying the same base is add the powers \[x^a \times x^b = x^{a + b}\] so for your question you need to use Numbers, Letters Powers Signs \[(-5mn^3)(4m^2n^2) = -5 \times 4m^{1 + 2}n^{3 + 2}\] just simplify it from here. hope it helps